OpenStudy (anonymous):

Larry Wants To Buy Some Carpeting For His Living Room. The Length Of The Room Is 2 Times The Width And The Total Area Of The Room Is 18 Square Meters. What Is The Length Of The Living Room

OpenStudy (anonymous):

@Misthorn

OpenStudy (mayankdevnani):

l=2*w

OpenStudy (mayankdevnani):

a=l*w

OpenStudy (mayankdevnani):

18=2w*w 18=2w^2

OpenStudy (mayankdevnani):

can you solve it @DamnImBad

OpenStudy (anonymous):

l =2w area =w*l =2w^2 18/2=w^2 9 = w^2 w= 3 l =2*3 =6

OpenStudy (mayankdevnani):

@surdawi don't give direct answers.

OpenStudy (anonymous):

When I Calulated It All Its Answer Was 6

OpenStudy (anonymous):

@DamnImBad next time put your work and we can check it for you

OpenStudy (mayankdevnani):

yaa you are correct

OpenStudy (mayankdevnani):

ok anything problem

OpenStudy (anonymous):

Thanks EveryOne ,

OpenStudy (mayankdevnani):

welcome